OpenStudy (andriod09):

I'm sure this is easy, but I haven't done this in about 4 years, need help solving \[4x\times5\]

OpenStudy (amistre64):

as is, there is no "solution" other than: the set of Real numbers

OpenStudy (andriod09):

i have to "Simplify the Following."

OpenStudy (amistre64):

oh, but its pretty simple as is :) multiply what you can multiply, and leave as is the things you cant

OpenStudy (andriod09):

But i was taught you can't multiply until you found the variable, and if i do that, i don't want to deal with the decimal

OpenStudy (amistre64):

simplification is just rewriting an expression; it has nothing to do with "solving for x" the variable is a placeholder is all

OpenStudy (amistre64):

4x*5 = 4*x*5 = 4*5*x = 5*4*x = 20*x = 20x

OpenStudy (amistre64):

4x = x+x+x+x 5 times, 4x = (x+x+x+x) + (x+x+x+x) + (x+x+x+x) + (x+x+x+x) + (x+x+x+x) = 20x

OpenStudy (andriod09):

Ah okay. that's simpler than I thought. :P Again, I haven't done this simple of math in a very long time. Thanks.

OpenStudy (amistre64):

youre welcome

OpenStudy (andriod09):

It would be negative 20x, if there was a negative number, right? it was \[4x\times-5\] was the real equation.